OpenStudy (anonymous):

write each equation in standard form then solve 21x^2 +5x-35=3x^2-4x

OpenStudy (anonymous):

? where is the question

OpenStudy (anonymous):

you can take this one @texaschic101 XD

OpenStudy (tkhunny):

Have you considered moving everything to the left-hand side, combining like terms, and then solving?

OpenStudy (anonymous):

so -3x^2-4x?

OpenStudy (anonymous):

or +4x i mean

OpenStudy (tkhunny):

Not close. Where did the "=" go?

OpenStudy (texaschic101):

21x^2 + 5x - 35 = 3x^2 - 4x 21x^2 - 3x^2 + 5x + 4x - 35 = 0 17x^2 + 9x - 35 = 0 now use the quadratic formula http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

ok but how did you go from 21x^2-3x^2+5x+4x-35=0 to 17x^2=9x-35=0?

OpenStudy (tkhunny):

21-3 = 18

OpenStudy (anonymous):

ok then i multiply 18 times -35?

OpenStudy (tkhunny):

Why would you do that? Factor or use the quadratic formula.

OpenStudy (anonymous):

or do i just do the x method

OpenStudy (tkhunny):

There are many names for many things. If that is what you mean for finding the factors, then yes. If that is what you mean for identifying your mutant powers, then no.

OpenStudy (anonymous):

I can't find the factors that add to 9 but multiply to -35

OpenStudy (tkhunny):

Super. Have you considered the quadratic formula? Maybe no such factors exist.

OpenStudy (anonymous):

i'm so confused

OpenStudy (tkhunny):

21x^2 +5x-35=3x^2-4x 18x^2 + 9x - 35 = 0 9^2 - 4(18)(-35) = 81 + 2520 = 2601 \(\sqrt{2601} = 51\) - And this is where we feel bad because it was a perfect square and we should have been able to factor it. Since we didn't, we have... \(x = \dfrac{9 \pm 51}{42}\) That should fill you in on how to factor it.

OpenStudy (anonymous):

Just confused me even more but nevermind i'll just try to ask my teacher for more help